40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 18 02:21:41 2013, Koji, Update, LSC, PRMI+Y arm ALS success! Screenshot-Untitled_Window1.pngPRMI_Yarm.pdf
    Reply  Thu Jul 18 10:47:21 2013, Jamie, Update, LSC, PRMI+Y arm ALS success! 
    Reply  Thu Jul 18 10:50:54 2013, Lisa, Update, LSC, PRMI+Y arm ALS success! 
    Reply  Thu Jul 18 15:55:31 2013, Jenne, Update, LSC, PRMI+Y arm ALS Sensing Matrices SensMatMeas_17July2013_PRMI_YarmOffResonance.pngSensMatMeas_17July2013_PRMI_YarmMidResonance.pngSensMatMeas_17July2013_PRMI_YarmOnResonance.png
       Reply  Thu Jul 18 16:30:08 2013, Koji, Update, LSC, PRMI+Y arm ALS Sensing Matrices 
    Reply  Fri Jul 19 12:02:18 2013, manasa, Update, LSC, PRMI+Y arm ALS success! 
Message ID: 8871     Entry time: Thu Jul 18 15:55:31 2013     In reply to: 8867     Reply to this: 8872
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI+Y arm ALS Sensing Matrices 

Last night, I took sensing matrix data at various different offsets for the Yarm.  The sensing matrices I measured were of the PRMI, while the Yarm was (a) Held off resonance, (b) Held at ~50% peak power, and (c) Held on resonance.

The dither lines were clear in the MICH and PRCL spectrum, so I think I'm driving hard enough, but something else seems funny, since clearly the REFL165 I and Q signals were not completely overlapping last night.  If they were, we wouldn't have been able to lock the PRMI using REFL 165 I&Q.

Anyhow, here's the data that was taken.  Data folder is ...../scipts/LSC/SensingMatrix/SensMatData/

Yarm off resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_012848.dat

SensMatMeas_17July2013_PRMI_YarmOffResonance.png

Yarm at ~50% resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_013937.dat

SensMatMeas_17July2013_PRMI_YarmMidResonance.png

Yarm on resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_013619.dat

SensMatMeas_17July2013_PRMI_YarmOnResonance.png

 

ELOG V3.1.3-